OpenStudy (anonymous):

Can someone help me solve some derivatives?pls i have solution pls check if it's correct and pls help me continue solving the others

OpenStudy (anonymous):

Sure, give me your answer and question.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Go ahead and post your answer/question and I'll try to explain it.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=m3im-iJdhv4

OpenStudy (anonymous):

go ahead and post your question i will try and explain it

OpenStudy (anonymous):

y=x/x^2-1 u=x v=x^2-1 du=1 dv=2x dy/dx=v*du/dx-u*dv/dx/v^2 =x^2-1(1)-x(2x)/(x^-12)^2 =(x^2-1)-2x^2

OpenStudy (anonymous):

/(x^2-1)^2

OpenStudy (anonymous):

what should i do to 2x^2?

OpenStudy (anonymous):

\[(x)/(x^2-1)\] is that the orginal function?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Oh dear... Partial derivatives. My apologies, but I'm in cal 1, so I dunno how to do those yet. :/

OpenStudy (anonymous):

the answer is x^2+1/(x^2-1)^2

OpenStudy (anonymous):

well, you could simply apply the quotient rule

OpenStudy (anonymous):

which you did do, except you used variables

OpenStudy (anonymous):

i'm applying it but i,m confused on the npart i stopped solving

OpenStudy (anonymous):

how did it become x^2+1?

OpenStudy (anonymous):

the numerator

OpenStudy (anonymous):

Lets do through this step by step, okay, except this time we wont use variable okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.twiddla.com/586786